Tom’s uncle owns a triangular piece of land. The perimeter fence that surrounds the land measures 378 yards. The shortest side is 30 yards longer than one-half of the longest side. The second longest side is 2 yards shorter than the longest side. What is the length of each side?
Answer by checkley77(12844)

.5X+30+X-2+X=378
2.5X=378-30+2
2.5X=350
X=350/2.5
X=140 IS THE LENGTH OF THE LONGEST SIDE.
.5*140+30=70+30=100 IS THE SHORTEST SIDE.
140-2=138 FOR THE MIDDLE LENGTH SIDE.
pROOF:
140+100+138=378
378=378